OpenStudy (anonymous):

WILL FAN AND MEDAL FOR ANSWER ! Adanna had a glass cylinder with a diameter of 23 cm and a height of 10 cm. She placed 240 marbles, each with a diameter of 1.2 cm, into the cylinder. She filled the rest of the cylinder with colored water. How many cubic centimeters of colored water did she use? Use 3.14 to approximate pi and round your answer to one decimal place.

OpenStudy (welshfella):

first find the volume of the cylinder = 3.14*r^2*h then find the volume of the 240 marbles volume of 1 marble = (4/3)*3.14* r^3

OpenStudy (welshfella):

volume of colored water = volume of cylinder - total volume of the240 marbles

OpenStudy (anonymous):

What would be the answer ?

OpenStudy (welshfella):

note :- radius r = half the diameter

OpenStudy (welshfella):

I can't tell you that - if you plug in the values in the question you can work it out as I explained

OpenStudy (chosenmatt):

we are here to guide you to the answer not to tell you the answer :) and make sure you understand what we are showing you :)

OpenStudy (igreen):

First find the volume of the cylinder. \(V = (3.14) (11.5)^2 (10)\) Simplify that @princess_hannahleigh

OpenStudy (welshfella):

Volume of cyl = 3.14 * 11.5^2 * 10 volume of 24 marbles = (4/3) *3.14* (0.6)^3

OpenStudy (anonymous):

\[V_{cylinder}=(240 \times V_{marble})+V_{water}\]

OpenStudy (anonymous):

\[V_{cylinder}=\pi r^2h\]\[V_{sphere}=\frac{4}{3}\pi r^3\]

OpenStudy (igreen):

\(V_{water} = V_{cylinder} - V_{marbles}\)

OpenStudy (welshfella):

oops Volume of the marbles is 240 times what I wrote!!

OpenStudy (anonymous):

Use the formulae everyone has stated and attempt the question

OpenStudy (igreen):

@princess_hannahleigh Multiply 3.14 * 11.5 * 11.5 * 10

OpenStudy (chosenmatt):

i gave you a medal @welshfella nice job on explaining :)

OpenStudy (anonymous):

4,152.65 @iGreen

OpenStudy (igreen):

Yes, now we find the volume of each marble and multiply by 240. diameter = 1.2 raidus = 1.2 / 2 = 0.6 Volume of a Sphere: \(V = \dfrac{4}{3} \pi r^3\) \(V = \dfrac{4}{3} (3.14)(0.6^3)\) Now multiply 4/3 * 3.14 * 0.6 * 0.6 * 0.6

OpenStudy (igreen):

Can you do that? @princess_hannahleigh

OpenStudy (anonymous):

0.90432

OpenStudy (igreen):

Yes, now multiply by 240

OpenStudy (anonymous):

217.0368

OpenStudy (igreen):

Yep, now subtract 4,152.65 - 217.0368

OpenStudy (anonymous):

3,935.6132

OpenStudy (igreen):

Yes, that's how much water was in the cup.

OpenStudy (igreen):

Well, it says to round it to one decimal place, so we round to 3,935.6

OpenStudy (anonymous):

So the answer is 3,935.6

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

Thanks so much!